OpenStudy (anonymous):

simplify sqrt 20b^3c^4

OpenStudy (gw2011):

square root (20b^3c^4)=square root (5)(4)(b^2)(b)(c^4)=2bc^2 (square root 5b)

OpenStudy (anonymous):

\[2bc ^{2} \sqrt{5b}\]